OpenStudy (anonymous):

Hi hero, if you have some time can you help me understand how to solve Quadratic Equations, please?

OpenStudy (anonymous):

\[-(-7_)\pm \sqrt{(-7)^{2}-4(3)(-5)} \over2(3)\]

OpenStudy (anonymous):

the next step is

OpenStudy (anonymous):

\[7\pm \sqrt{109}\over6\]

OpenStudy (turingtest):

looks good so far :)

OpenStudy (anonymous):

i don't understand the result 109

OpenStudy (anonymous):

for instance: i know 7 sq = 49

OpenStudy (anonymous):

4 * 3 = 12

OpenStudy (anonymous):

4*-5=-20

OpenStudy (anonymous):

but they don't add up to 109

OpenStudy (turingtest):

you just can't simplify \[\sqrt{109 }\]so just leave it like that

OpenStudy (anonymous):

this is sample equation

OpenStudy (turingtest):

oh I see what you mean...

OpenStudy (anonymous):

lets say the 2nd equation doens't exist: how to I arrive to the results of the 2nd equation

OpenStudy (turingtest):

you have that under the radical (which is called the discriminant by the way) is \[(-7)^2-4(3)(-5)\]right? and you just don't see how that equals 109?

OpenStudy (anonymous):

yes!

OpenStudy (turingtest):

\[(-7)^2-4(3)(-5)=49+4(3)(5)=49+12(5)=49+60=109\]you maybe just missed the negative in the -5 makes the terms add I'm guessing....

OpenStudy (anonymous):

tytytytytyty i needed to see it just like that. tytytytyt

OpenStudy (turingtest):

anytime :)